If f(x) = 3x-4 and g(x) = -x+5, find the equations for f)g and g)f.
Confusing.. I am blonde.. haha.. I do need some help though. Thankyou Shelly.

Answer by scott8148(6628)

me too...haha...

f(g(x))=3(g(x))-4=3(-x+5)-4=-3x+15-4=-3x+11

g(f(x))=-(f(x))+5=-(3x-4)+5=-3x+4+5=-3x+9